2x∧2+6x+7=2 completeing the square formula
Answer by stanbon(75887)

2x∧2+6x+7=2 completing the square formula
--------------------
2(x^2+3x+?) = 2-7 + 2*?
---
2(x^2+3x+(3/2)^2) = 2-7+2(3/2)^2
2(x+(3/2))^2 = -5 + (9/2)
2(x+(3/2))^2 = -1/2
===============================
Cheers,
Stan H.